# PATCH! Psychometrics-AssisTed benCHmarking of Large Language Models: A Case Study of Proficiency in 8th Grade Mathematics

### v2

@@ 3.3 LLM Benchmark Development
Developing LLM benchmarks follows a similar yet different process. Take GSM8K (Cobbe et al., 2021) as an example. According to the GSM8k paper, the authors started by specifying the need for a large, high quality mathematics test at grade school level that is of moderate difficulty for LLMs (Step 1). The implied construct (i.e., PGSM) is, however, not explicitly linked to any specific curriculum.
Then, the overall planning is made (Step 2): The number of items should be in the thousands; the items will be curated by crowd workers; agreement and error analysis will be used to investigate the quality of the dataset; GPT-3 will be used to benchmark the dataset and verify the difficulty of the dataset.
In Step 3, where dataset development [^8] Note that we use the term “dataset development” here, contrasting “item development” in psychometrics, because of LLM benchmarks’ typical emphasis on large and multiple datasets rather than concrete test items. takes place, often one of the two strategies is used: either collect items from existing datasets and other sources and compile them into a new dataset, or, like in GSM8K, create own items from scratch (with annotations). The latter is usually an iterative procedure consisting of four parts: creating instructions (and possibly a user interface) for item generation and/or annotation; conduct a (pilot) study to collect the items and/or annotations; check annotator agreement; and assess errors associated with the items or annotations. This step is iterated until a sufficient number of items and datasets is reached while meeting desired quality standards (e.g., high annotator agreement, low error rate). In total, GSM8K includes 8,500 items with solutions, with identified annotator disagreements resolved and a less than 2% error rate.
In Step 4, the generated items make up the final dataset, typically split into training, evaluation and testing partitions. In Step 5, the final selection of LLMs is made and evaluated on the dataset. Finally, in Step 6, the benchmark gets released, which typically consists of the dataset as well as its documentation (often a research paper) and benchmarking results.
@@
While sharing similarity with test development in psychometrics, benchmark development for LLMs falls short on four aspects. First, the construct of interest is often under-specified, leading to a mismatch between the intended construct and what the dataset actually measures. Take GSM8K as an example: While the dataset is intended to measure proficiency in grade school mathematics, the target grade level(s) are unclear and it only focuses on one content domain (algebra), missing other relevant ones like geometry and data. This is likely the result of not using established mathematics curricula and domain experts to develop test items.
@@
Third, besides agreement and error analysis, LLM benchmarks can benefit from psychometric analysis of test items, (i.e., checking item discriminativeness and difficulty, as well as the factor structure of the items). While this is not yet the norm, there have been promising attempts (see Section 2).
Lastly, the released benchmark often does not contain sufficient details about all the steps involved in creating the benchmark. For instance, the GSM8K paper does not present the instructions for item creation and annotation, the results from the pilot study, the agreement statistics, or annotator characteristics, all of which are important for external researchers to independently validate the quality of the benchmark.
@@ 4 PATCH: Psychometrics-AssisTed benCHmarking of LLMs
Figure 1 illustrates PATCH, our conceptualisation of a psychometrics-assisted framework for benchmarking LLMs. [^9] PATCH is partly inspired by the Hexagon Framework of scientific measurements proposed by Mari et al. (2023). According to PATCH, the first step is to define the construct of interest (e.g., proficiency in 8th grade mathematics).
@@
The second step is to look for an existing validated psychometric test that measures this property; alternatively, a test can be developed from scratch, following the procedures described in Section 3.2, which likely requires collaboration with experienced psychometricians. The term “validated” means that the test has been tested on a representative sample of the target population of human test takers and fulfils several psychometric quality requirements (e.g., discriminative items that are well distributed across different difficulty levels; showing high reliability (e.g., high internal consistency) and validity (e.g., the test’s factor structure matches the construct definition)).
Figure 1: PATCH: A Psychometrics-AssisTed framework for benCHmarking LLMs.
Next (Step 3$\rightarrow$4), we use the items in the validated psychometric test to construct prompts for the LLMs under evaluation and then sample responses. A response typically consists of a task description, an explanation and an answer (key). Therefore, in Step 4$\rightarrow$5, we extract the answer (key) for each item’s response, then grade it to obtain item scores (Step 5$\rightarrow$6).
@@
For Step 2$\rightarrow$7, the responses of human test takers (and of LLMs, if a sufficient number of LLMs are involved) can be used to estimate IRT item parameters and subsequently the latent proficiency scores for each test taker (human or LLM) with uncertainty estimates. Multiple IRT models are often used because of the use of different types of test items. These latent scores are typically standardised $z$-scores (i.e., mean of 00 and standard deviation of $1$), which can optionally go through further normalisation (e.g., re-scaling to a mean of $500$ and a standard deviation of $100$) (Step 6$\rightarrow$7). These final proficiency scores can be used for comparison with other models and populations.
@@
It can be said that the heart of PATCH is a validated psychometric test, which not only provides the basis for accurate measurement of a model’s capability of interest but also facilitates comparison between LLMs and human test takers. Unfortunately, developing such a test can be a long and expensive process; utilising existing tests can be a shortcut, which, however, should satisfy three requirements: clear human population reference; test items available (released); human responses and/or item parameter estimates available. The second and third requirement are in practice difficult to meet, as many test institutes do not make their test items public due to commercial interests (e.g., SAT) or the need to measure trends over time (e.g., PISA). Collaboration with test institutes would alleviate this problem.
@@
To the best of our knowledge, when it comes to academic proficiency tests, only TIMSS and PIRLS tests from certain years can be readily used for PATCH-based LLM benchmarking. TIMSS measures proficiency in grade school mathematics and science (4th grade, 8th grade, and final year of secondary school), while PIRLS assesses reading comprehension in 9/10-year-olds. Both TIMSS and PIRLS are administered in a large number of countries and regions with representative student samples, enabling country/region-level comparisons. In the following section, we demonstrate PATCH by measuring GPT-4 and Gemini’s proficiency in 8th grade mathematics, using the latest available data from TIMSS 2011.
5 Demonstration: Measuring LLM Proficiency in 8th Grade Mathematics
@@ 5.1 Data: TIMSS 2011 8th Grade Mathematics
56 countries/regions participated in TIMSS 2011, with typically a random sample of about 150 schools in each country/region and a random sample of about 4,000 students from these schools. These sample sizes are determined on the basis of a .$\leq.035$ standard error for each country’s mean proficiency estimate. The use of random sampling makes unbiased proficiency estimates possible at the population level. TIMSS 2011 has released a publicly available database [^10] https://timssandpirls.bc.edu/timss2011/international-database.html, of which three components are relevant to our study:
@@
The TIMSS 2011 study has released 88 mathematics test items, 48 of which are multiple choice, 30 open-ended items scored as either incorrect or correct, and 10 open-ended items scored as either incorrect, partially correct, or correct. These items assess four content domains representative of 8th grade mathematics curriculum (agreed upon by experts from participating countries/regions): number, algebra, geometry, data and chance. Within each domain, items are designed to cover various subtopics (e.g., decimals, functions, patterns) and three cognitive domains: knowing, applying and reasoning. These test items are only available in a PDF file that can be downloaded from the NCES website, which includes also scoring instructions. [^11] https://nces.ed.gov/timss/pdf/TIMSS2011_G8_Math.pdf To extracting them into a format compatible with LLMs, we used OCR tools to extract as much textual information as possible, converted mathematical objects (e.g., numbers, symbols, equations, tables) into LaTeX format (following earlier benchmarks like MATH (Hendrycks et al., 2021)) and figures into JPEG format. See Appendix A.1 for examples. We have released this LLM-compatible version of test items, as well as an eighth grade science test dataset from TIMSS 2011, an advanced secondary school mathematics test dataset from TIMSS 2008, and an advanced secondary school physics test dataset from TIMSS 2008121212Available via https://zenodo.org/records/12531906.
@@
The second part of the dataset concerns the specific IRT model used for each test item and the estimated item parameters (e.g., discriminativeness, difficulty), which can be used to reconstruct the IRT models for estimating proficiency scores.
@@
Lastly, responses of the sampled students (about 4,000 on average per country/region) to each test item and their proficiency estimates have also been made available, allowing us to construct proficiency score distributions for each country and region.
5.2 LLMs: GPT-4 with Vision and Gemini-Pro-Vision
Considering that more than 1/3 of the test items contain visual elements, we chose two vision language models: GPT-4 with Vision (GPT-4V) and Gemini-Pro-Vision, using the respective APIs. We are aware of other LLMs with vision capabilities. However, our goal is to showcase PATCH instead of benchmarking all relevant LLMs.
A major concern in using these closed-source LLMs is data contamination, which is difficulty to check due to inaccessible (information about) training data. However, as our focus is on demonstrating the PATCH framework, data contamination is less worrying. Furthermore, data contamination is still unlikely for four reasons. First, these test items are copyrighted, forbidding commercial use. Second, the test items are hard to extract from the PDF file. Third, to the best of our knowledge, these test items do not exist in current LLM mathematics benchmarks. Fourth, we ask GPT-4V and Gemini-Pro-Vision to explain or provide solutions to the test items’ IDs (available in the PDF file). Both failed to recognise these specific test IDs.
@@ 5.3 Prompts and Temperature
We design two separate prompts for each test item: the system message and the user message. We design the system message according to the prompt engineering guide by OpenAI, utilising chain-of-thought and step-by-step instructions on how to respond to the user message (i.e., with a classification of question type, an explanation and an answer (key)). [^13] https://platform.openai.com/docs/guides/prompt-engineering The system message is the same for all test items (see Appendix A.2). Furthermore, to account for LLMs’ sensitivity to slight variations in prompts (e.g., Sclar et al., 2024; Loya et al., 2023), we generate 10 additional variants of the system prompt with slight perturbations (e.g., lowercase a heading, vary the order of unordered bullet points).
The user message is item-specific, containing both the item’s textual description and the associated image(s) in base 64 encoded format. See Appendix A.1 for examples. [^14] While we are aware of other prompt engineering techniques, such as few-shot prompting and self-consistency, we did not experiment with them, as our focus is on demonstrating the use of PATCH.
@@
5.4 Response Scoring and Proficiency Estimation
We manually examine the sampled responses from GPT-4V and Gemini-V and score them following the official scoring rubrics of TIMSS 2011. Then, for multiple choice items, we apply the 3PL model (Equation 1); for open-ended items, we apply the GPC model (Equation 2) if partially correct response is admissible, otherwise the 2PL model. We use maximum likelihood to obtain unbiased estimates of model proficiency scores ($\theta$) with the mirt package in R (Chalmers, 2012). This results in 11 $\theta$ estimates per model due to the use of 11 system message variants. We then use inverse variance weighting (Marín-Martínez & Sánchez-Meca, 2010) to combine these estimates. Inverse variance weighting gives more weight to estimates that are more precise (i.e., having lower variance) and less weight to those that are less precise (i.e., having higher variance). This way, we obtain a more accurate overall $\theta$ estimate and its 95% confidence interval (CI) for each model.
@@ 5.5 Results
Figure 2: Distribution of proficiency estimates for GPT-4V, Gemini-Vision-Pro and selected participating countries/regions of the TIMSS 2011 8th grade mathematics test. Left figure (A) shows the proficiency estimates based on the percentages of correct responses. Right figure (B) shows the IRT-based proficiency estimates. The middle vertical line in each box plot represents the weighted mean proficiency score, with the error bars indicating its 95% confidence interval. The borders of each box indicate the range of the middle 50% of all values, with the two whiskers indicating the 5th and 95th percentiles.
Figure 2 shows the proficiency score distribution and ranking of 15 selected participating countries and regions, GPT-4V and Gemini-Pro-Vision. Only 15 countries are shown here to save space. The complete figures can be found in Appendix C. The proficiency scores (x-axis) on the left panel are percentages of correct responses, which is the default approach in current LLM benchmarking; the proficiency estimates on the right panel are based on IRT. We make three observations. First, regardless of the method of proficiency estimation, GPT-4V has the overall best performance relative to Gemini-Pro-Vision and the average proficiency of 8th grade students of each participating country/region. Second, the method of proficiency estimation affects the ranking results. For instance, while Chinese Taipei is ranked 3rd on the left, it is ranked 4th on the right; Gemini-Pro-Vision is ranked 8th on the left, but ranked 7th on the right. Similarly, while Hungary is ranked 11th on the left, it drops to the 16th place on the right. Third, the method of proficiency estimation affects the estimated 95% CIs, which are usually wider when IRT is used (as it accounts for both item and test taker variances). Notably, while on the left panel the CI of GPT-4V does not overlap with the second best, South Korea, indicating a statistically significant difference, they overlap on the right panel, suggesting otherwise. This finding shows that the adoption of PATCH is likely going to make a difference to LLM benchmark results.
6 Conclusions

### v3

@@ 3.3 LLM Benchmark Development
In this section, we focus on the right panel of Table 1: the process of LLM benchmark development, contrast it with test development in psychometrics and thereby highlight the potential benefits a psychometrics-based approach can introduction to LLM benchmarks.
The process of developing LLM benchmarks is similar to test development in psychometrics. However, there are significant differences. To illustrate this, we take GSM8K Cobbe et al. (2021) as an example, where we try our best to recreate the process of developing GSM8k based on the published dataset paper and map the specific steps to the six steps described in the right panel of Table 1.
First, the authors of GSM8K likely started by specifying the need for a large, high quality mathematics test at grade school level and of moderate difficulty for LLMs (Step 1). However, they did not explictly link the construct (i.e., PGSM) to any specific curriculum. Then, the authors made overall planning for the benchmark development (Step 2). For example, the number of items should be in the thousands; the crowd workers would curate the benchmark items; the authors would use agreement and error analysis to investigate the quality of the dataset; GPT-3 will be used to benchmark the dataset and verify dataset difficulty.
In Step 3, namely dataset development [^6] Note that we use the term “dataset development” here, contrasting “item development” in psychometrics, because of LLM benchmarks’ typical emphasis on large and multiple datasets rather than concrete test items., often one of the two strategies is used: either collect items from existing datasets and other sources and compile them into a new dataset, or, create own items from scratch (with annotations). The authors of GSM8K followed the latter approach, an iterative procedure consisting of four parts: creating instructions (and possibly a user interface) for item generation and/or annotation; conducting a (pilot) study to collect the items and/or annotations; check annotator agreement; and assessing errors associated with the items or annotations. This step is iterated until a sufficient number of items and datasets are reached while meeting desired quality standards (e.g., high annotator agreement, low error rate). In total, GSM8K includes 8,500 items with solutions, with identified annotator disagreements resolved and a less than 2% error rate.
In Step 4, the GSM8K authors compiled the final dataset from the crowdsourced items with training, evaluation and testing partitions. In Step 5, the GSM8k authors evaluated selected LLMs (i.e., GPT-3) on the dataset. Finally, in Step 6, the authors released the benchmark, which consists of the dataset as well as its documentation (a research paper) and benchmarking results.
@@
While sharing similarity with test development in psychometrics, current benchmark development for LLMs falls short on four aspects. First, the construct of interest is often under-specified, leading to a mismatch between the intended construct and what the dataset actually measures. Again, take GSM8K as an example: While the dataset is intended to measure proficiency in grade school mathematics, the target grade level(s) are unclear and it only focuses on one content domain (algebra), missing other relevant ones like geometry and data. This is likely the result of not using established mathematics curricula and domain experts to develop test items.
@@
Third, besides agreement and error analysis, LLM benchmarks can benefit from psychometric analysis of test items, (i.e., checking item discrimination and difficulty, as well as the factor structure of the items). While this is not yet the norm, there have been promising attempts (see Section 2).
Lastly, the released benchmark often does not contain sufficient details about the process of benchmark creation. For instance, the GSM8K paper does not report instructions for item generation and annotation, results of the pilot study, agreement statistics, or annotator characteristics, all of which are important for external researchers to independently verify the quality of the benchmark.
@@ 4 PATCH: Psychometrics-AssisTed benCHmarking of LLMs
Figure 1 illustrates PATCH, our conceptualisation of a Psychometrics-AssisTed framework for benCHmarking LLMs against human populations. Each box represents a different research artifact, while each arrow applies an action to a source artifact and produces a subsequent target artifact.
@@
Under PATCH, the first step is for researchers to define the construct of interest (e.g., proficiency in 8th grade mathematics), the specific LLM under examination, as well as the reference human population for comparison (e.g., 8th graders in Germany). Then, researchers look for an existing validated [^7] The term “validated” means that the test has been (pre)tested on a representative sample of the target population of (human) test takers and fulfils psychometric quality requirements, such as sufficiently many discriminative items well distributed across different difficulty levels, and showing high reliability (e.g., high internal consistency) and validity (e.g., the sizes and directions of the empirical correlations among test items match theoretical expectations). psychometric test measuring the specified construct; alternatively, a test can be developed from scratch by following the procedures described in Section 3.2, which likely requires collaboration with experienced psychometricians.
Next, researchers use the items from the validated psychometric test to construct prompts for the LLM under evaluation and sample responses from the LLM. Similarly, researchers administer the psychometric test to a representative sample of the reference human population, and collect their responses. These human responses are then used to train and validate the IRT model(s) that match the type(s) of items in the psychometric test.
@@
The resulting IRT model(s) will be fitted to the responses from the LLM and the humans, and will subsequently estimate each test taker’s latent proficiency score—whether human or LLM—along with uncertainty estimates. These final proficiency scores [^8] These latent proficiency scores are typically standardised $z$-scores (i.e., mean of 00 and standard deviation of $1$), which sometimes go through further normalisation (e.g., re-scaling to a mean of $500$ and a standard deviation of $100$). enable valid comparison between the LLM and the reference human population.
@@
At the heart of PATCH lies the psychometric test, which not only provides the basis for accurate measurement of the construct (i.e., capability of interest) but also enables valid comparison between LLMs and human test takers. Unfortunately, developing such a test can be a long and expensive process; utilising existing tests can be a shortcut, which should satisfy three conditions: 1) clear human population reference; 2) test items available in the public domain; 3) human responses and/or item parameter estimates available. The second and third are in practice difficult to meet, as many test institutes do not make their test items public due to commercial interests (e.g., SAT) or the need to measure trends over time (e.g., PISA). Collaboration with test institutes would alleviate this problem and additionally mitigate the likely data contamination issue with many public benchmarks.
@@
To the best of our knowledge, among academic proficiency tests, only TIMSS and PIRLS tests from certain years can be readily used for PATCH-based LLM benchmarking without formal collaboration with test institutes. TIMSS measures proficiency in grade school mathematics and science (4th grade, 8th grade, and final year of secondary school), while PIRLS assesses reading comprehension in 9/10-year-olds. Both TIMSS and PIRLS are administered in a large number of geographical regions with representative student samples, enabling population-level comparisons. In the following section, we demonstrate PATCH by measuring several LLMs’ proficiency in 8th grade mathematics, using the latest available data from TIMSS 2011.
5 Demonstration: Leveraging IRT in PATCH to Measure LLM Proficiency in 8th Grade Mathematics
In this section, we conduct an experiment to demonstrate the benefits and differences IRT modelling can make under our PATCH framework, compared to the traditional benchmark scoring approach. Note that we focus on the IRT part of PATCH instead of also on test development, because we lack the resources to develop our own valid test, and that there is no existing LLM benchmark that has the same measurement target as TIMSS 2011 (i.e., 8th grade mathematics proficiency) to enable comparison in terms of test development. Nevertheless, we hope our detailed comparison between LLM benchmark devleopment and psychomtric test development in Section 3.3 suffices to fill in this gap.
@@ 5.1 Data: TIMSS 2011 8th Grade Mathematics
56 geographical regions participated in TIMSS 2011, with typically a random sample of about 150 schools in each region and a random sample of about 4,000 students from these schools. These sample sizes are determined on the basis of a .$\leq.035$ standard error for each region’s mean proficiency estimate. The use of random sampling makes unbiased proficiency estimates possible at the population level. TIMSS 2011 has released a publicly available database [^9] https://timssandpirls.bc.edu/timss2011/international-database.html, of which three components are relevant to our study:
@@
The TIMSS 2011 study has released 88 mathematics test items, 48 of which are multiple choice, 30 open-ended items scored as either incorrect or correct, and 10 open-ended items scored as either incorrect, partially correct, or correct. These items assess four content domains representative of 8th grade mathematics curriculum (agreed upon by experts from participating regions): number, algebra, geometry, data and chance. Within each domain, items are designed to cover various subtopics (e.g., decimals, functions, patterns) and three cognitive domains: knowing, applying and reasoning. These test items are only available in a PDF file that can be downloaded from the NCES website, which includes also scoring instructions. [^10] https://nces.ed.gov/timss/pdf/TIMSS2011_G8_Math.pdf To extract them into a format compatible with LLMs, we used OCR tools to extract as much textual information as possible, converted mathematical objects (e.g., numbers, symbols, equations, tables) into LaTeX format (following earlier benchmarks like MATH) Hendrycks et al. (2021) and figures into JPEG format. See Appendix A.1 for examples. We have released this LLM-compatible version of test items, as well as an eighth grade science test dataset from TIMSS 2011, an advanced secondary school mathematics test dataset from TIMSS 2008, and an advanced secondary school physics test dataset from TIMSS 2008. See Appendix C for details.
@@
The TIMSS 2011 database also specifies the IRT model used for each test item and contains the item parameter estimates (e.g., discrimination, difficulty), which we use to reconstruct the final IRT model for proficiency estimation and verification.
@@
Lastly, responses of the sampled students to each test item and their proficiency estimates are also available, allowing us to construct proficiency score distributions for each region.
5.2 LLMs: GPT-4, Gemini-Pro and Qwen with Vision Capability
Considering that more than 1/3 of the test items contain visual elements, we selected four competitive vision language models: GPT-4 with Vision (GPT-4V), Gemini-Pro-Vision, as well as the open-source Qwen-VL-Plus and Qwen-VL-Max Bai et al. (2023). There are more LLMs with vision capability. However, our goal is to showcase PATCH, not to benchmark as many LLMs as possible.
A major concern in using these LLMs is data contamination, which is difficulty to check due to inaccessible (information about) training data. However, as our focus is on demonstrating the PATCH framework, data contamination is less worrying. Furthermore, data contamination is still unlikely for four reasons. First, these test items are copyrighted, forbidding commercial use. Second, the test items are hard to extract from the source PDF. Third, to the best of our knowledge, these test items do not exist in current LLM mathematics benchmarks. Fourth, we prompted the selected LLMs to explain or provide solutions to the test items’ IDs (available in the source PDF). All failed to recognise these specific test IDs.
@@ 5.3 Prompts and Temperature
We design two separate prompts for each test item: the system message and the user message. We design the system message according to the prompt engineering guide by OpenAI, utilising chain-of-thought and step-by-step instructions on how to respond to the user message (i.e., with a classification of question type, an explanation and an answer (key)). [^11] https://platform.openai.com/docs/guides/prompt-engineering The system message is the same for all test items (see Appendix A.2). Furthermore, to account for LLMs’ sensitivity to slight variations in prompts Sclar et al. (2024); Loya et al. (2023), we generate 10 additional variants of the system prompt with slight perturbations (e.g., lowercase a heading, vary the order of unordered bullet points).
The user message is item-specific, containing both the item’s textual description and the associated image(s) in base 64 encoded format. See Appendix A.1 for examples. [^12] We are aware of other prompt engineering techniques like few-shot prompting and self-consistency. We did not experiment with them, as our focus is on demonstrating PATCH.
@@
Figure 2: Distribution of proficiency estimates for GPT-4V, Gemini-Vision-Pro, Qwen-VL-Plus, Qwen-VL-Max and selected participating regions of the TIMSS 2011 8th grade mathematics test. Left figure (A) shows the proficiency estimates based on the percentages of correct responses. Right figure (B) shows the IRT-based proficiency estimates. The middle vertical line in each box plot represents the weighted mean proficiency score, with the error bars indicating its 95% confidence interval. The borders of each box indicate the range of the middle 50% of all values, with the two whiskers indicating the 5th and 95th percentiles. Note that we adhere to the official naming conventions of TIMSS 2011 when reporting the names of participating regions, with no intent to offend anyone.
5.4 Scoring and Proficiency Estimation
We manually scored the sampled responses from the LLMs following the official scoring rubrics of TIMSS 2011. Then, for multiple choice items, we apply the 3PL model (Equation 3.1); for open-ended items, we apply the GPC model (Equation 2) if partially correct response is admissible, otherwise the 2PL model. We use maximum likelihood to obtain unbiased estimates of model proficiency scores ($\theta$) with the mirt package in R Chalmers (2012). This results in 11 $\theta$ estimates per model corresponding to 11 system message variants. We then use inverse variance weighting Marín-Martínez and Sánchez-Meca (2010) to combine these estimates. Inverse variance weighting gives more weight to estimates that are more precise (i.e., having lower variance) and less weight to those that are less precise (i.e., having higher variance). This way, we obtain a more accurate overall $\theta$ estimate and its 95% confidence interval (CI) for each model. This further allows us to visually assess statistical significance by checking for overlap between CIs: for two independent samples, non-overlapping intervals suggest significance at $\alpha=0.01$; slight overlap may still imply significance at $\alpha=0.05$; and substantial overlap indicates non-significance at $\alpha=0.05$.
@@ 5.5 Results
Figure 2 shows the proficiency score distribution and ranking of the top 15 performing participating regions, as well as GPT-4V, Gemini-Pro-Vision, Qwen-VL-Plus and Qwen-VL-Max. The complete figures can be found in Appendix E. The proficiency scores (x-axis) on the left panel are percentages of correct responses, corresponding to the default approach in current LLM benchmarking; the proficiency estimates on the right panel are based on IRT. We make four observations:
First, regardless of the method of proficiency estimation, the proficiency estimates show that GPT-4V has the overall best performance relative to the other models and 8th grade students of each participating region.
Second, still looking at only the proficiency estimates, the method of proficiency estimation affects the ranking results. For instance, while Chinese Taipei is ranked 3rd on the left, it is ranked 4th on the right; Gemini-Pro-Vision is ranked 8th on the left, but ranked 7th on the right. Similarly, while Hungary is ranked 11th on the left, it drops to the 16th place on the right.
Based on the first and second observations, one might argue that the overall rankings are similar (despite some large deviations such as Hungary) and therefore, IRT does not make a real difference in benchmarking LLMs and human populations. However, the similarity between the results of these two estimation approaches is to be expected, as the TIMSS items are already validated and thus of high-quality (i.e., good measurement properties, of various difficulty levels, high discrimination), rendering the use of IRT-based proficiency estimates not necessarily substantially different from using simple aggregate scores. Had the items been inconsistent in terms of measurement quality, more notable differences would have been observed. Furthermore, our next observation, which focuses on uncertainty estimates, which are necessary on top of proficiency estimates for making claims about performance difference, lends support to our claimed importance of IRT.
Third, the method of proficiency estimation affects the estimated 95% CIs of human populations, which are usually wider when IRT is used. Notably, while on the left panel the CI of GPT-4V does not overlap with the second best (Rep. of Korea), indicating a statistically significant difference (t(10.26) = 3.19, p $<$ 0.01), they overlap on the right panel, suggesting otherwise (t(10.40) = 1.25, p = .24). This means that based on traditional proficiency estimation, GPT-4V performs significantly better than all the human groups, suggesting super-human performance. In contrast, when IRT is used, GPT-4V shares the same rank with Rep. of Korea, Singapore and Chinese Taipei, rejecting super-human performance.
Fourth, on the right panel, the CIs of the LLMs’ proficiency estimates are generally narrower than their counterparts on the left panel, indicating that using IRT leads to more precise proficiency estimates for LLMs, a further advantage of IRT.
These findings show that the adoption of IRT with PATCH is likely to make a difference to LLM benchmark results, especially in contrast with human performances.
6 Conclusion
In this paper, we propose PATCH, a psychometrics-inspired framework to address current limitations of LLM benchmarks, including questionable measurement quality, lack of quality assessment on the item level and unwarranted comparison between humans and LLMs. We demonstrate the IRT part of PATCH with an 8th grade mathematics proficiency test and show evaluation outcomes that diverge from those based on existing benchmarking practices, especially when comparison with human test takers is made. This underscores the potential of PATCH to reshape the LLM benchmarking landscape. Furthermore, we release 4 datasets that meet the requirements of PATCH, supporting the measurement of LLM proficiency in grade school math and science and its comparison with human performance. We hope to bring the LLM research community a step forward towards more scientific benchmarking and inspire more research in this direction. We also encourage researchers to collaborate with test institutes when developing new benchmarks, especially those that aim to measure cognitive capabilities.
Limitations